OpenStudy (anonymous):

8156.64mm^3/490.62mm^2 =16.62mm^2 or mm^3?

OpenStudy (phi):

You can treat units just as if they were variables. That is mm^3/mm^2 = mm (Just like x^3/x^2 = x). Use the rule of exponents: when dividing, subtract: 3-2= 1 Another example: if you were doing 10 mm times 10 mm you would get 10x10 mm x mm or 100 mm^2

OpenStudy (anonymous):

so if i was to turn it in what would i put?